The sum of three consecutive integers is 48. Find the integers

Answer by mananth(16946)


The sum of three consecutive integers is 48. Find the integers
n, n+1, n+2 be the integers
n+n+1+n+2=48
3n+3=48
3n = 48-3
3n=45
n= 15
n+1=16
n+2=17
15,16,17 are the integers